# Experimental Exploration: Investigating Cooperative Interaction Behavior Between Humans and Large Language Model Agents

### v1

@@ Abstract.
With the rise of large language models (LLMs), AI agents as autonomous decision-makers present significant opportunities and challenges for human-AI cooperation. While many studies have explored human cooperation with AI as tools, the role of LLM-augmented autonomous agents in competitive-cooperative interactions remains under-examined. This study investigates human cooperative behavior by engaging 30 participants who interacted with LLM agents exhibiting different characteristics (purported human, purported rule-based AI agent, and LLM agent) in repeated Prisoner’s Dilemma games. Findings show significant differences in cooperative behavior based on the agents’ purported characteristics and the interaction effect of participants’ genders and purported characteristics. We also analyzed human response patterns, including game completion time, proactive favorable behavior, and acceptance of repair efforts. These insights offer a new perspective on human interactions with LLM agents in competitive cooperation contexts, such as virtual avatars or future physical entities. The study underscores the importance of understanding human biases toward AI agents and how observed behaviors can influence future human-AI cooperation dynamics.
@@ 1. Introduction
The investigation into human decision-making processes within competitive and cooperative contexts has long been a cornerstone of behavioral economics, psychology, and neuroscience (Rand et al., 2015). This multidisciplinary research provides a nuanced understanding of the complexities involved in human decision-making. Over the past five years, the field of artificial intelligence has experienced significant growth, largely driven by the rapid advancement and deployment of large language models (LLMs) (Wu et al., 2023). These models are increasingly integrated into human society, influencing its development by performing tasks, interacting with humans, and making autonomous decisions (Liu et al., 2023). Historically, rule-based AI resolved specific problems through strictly defined logic by programmers and human experts (van der Waa et al., 2021). However, contemporary models now possess generalization capabilities, addressing issues not explicitly predefined. This shift exacerbates challenges with black-box models, making decision processes less transparent and harder to interpret (Nguyen et al., 2023). Increasing research highlights the effectiveness of LLMs in simulating human cognition and cooperative behavior (Ke et al., 2024; Abbasiantaeb et al., 2024), suggesting that the realization of concepts from works like “Detroit: Become Human [^1] https://www.quanticdream.com/en/detroit-become-human” is more imminent than anticipated (Dehnert and Leach, 2021; Strachan et al., 2024). Therefore, it is crucial to address the dynamics of cooperation between humans and AI agents, particularly as AI agents function as independent decision-makers interacting with humans in both competitive and cooperative scenarios (Yang et al., 2024a).
Existing research on human-AI interaction has primarily focused on AI serving as tools to augment human capabilities under human supervision (Xu et al., 2021; Amershi et al., 2019; Cabrera et al., 2023). However, with the rapid advancement of large language models (LLMs), the ability of AI agents to autonomously execute specific tasks has significantly improved (Luo et al., 2024), potentially even displacing certain job positions (De Cremer and Kasparov, 2021; Wang et al., 2024). Consequently, there is an urgent need to investigate how the AI agents influence human cooperative behavior in competitive scenarios between humans and AI agents.
To elucidate this issue, we used the Prisoners’ Dilemma as the experiment background to explores human cooperative behavior when they face to the different purported characteristic LLM agents. Our study aimed to investigate human with different gender cooperative behavior in the presence of AI agent competitors with varying characteristics and to evaluate how individuals respond to presumed competitors in cooperative scenarios. By experimentally examining three proposed hypotheses:
  • H1: The purported characteristics of LLM agents will significantly influence the cooperative behavior patterns of human participants in competitive environments.
  • H2: Participant gender will significantly affect the cooperative behavior patterns of individuals in competitive environments during interactions with LLM agents of differing purported characteristics.
  • H3: There is a significant interaction effect between the purported characteristics of LLM agents and the gender of human participants on the cooperative behavior patterns of human participants in competitive environments.
Finally, we found that participants exhibited a higher cooperation rate, more proactive favourable actions, and greater acceptance of repair efforts following betrayal when the AI agent was purported as human characteristics. Behavioral patterns, including the degree of proactive favourable actions and acceptance of repair efforts, varied among participants of different genders based on the AI agent’s purported characteristics. Moreover, participants had higher expectations of cooperation from LLM agents, but these elevated expectations also led to greater disappointment when the agents failed to meet them.
These findings provide empirical evidence on the vital factors influencing cooperative behavior between humans and AI agents in competitive environments, offering valuable insights for guiding developers in designing future AI agents. Additionally, this study contributes practical experience to enhance human-AI collaboration in the context of CSCW (Computer-Supported Cooperative Work) and to inform the development of AI agents as independent decision-makers in the future.
2. Background
In this section, we review the pertinent literature related to our research questions and elucidate how it underpins our investigation. We begin by examining prior studies on interaction behaviors between humans and AI agents within cooperative scenarios, analyzing their respective impacts. Subsequently, we explore competitive and collaborative environments, evaluating the metrics used to assess competitive decision-making. Finally, we investigate the role of AI characteristics in human-AI interactions and the influence of human participants’ gender on cooperative behaviors as documented in previous research.
2.1. Impact of AI Agents on Human Collaborative Behavior
@@
Large Language Models (LLMs) represent a significant advancement in AI, characterized by rapid response times, high adaptability, intelligent perception, and versatile communication capabilities (Chang et al., 2024). Through advanced natural language processing, LLM-driven agents can comprehend and generate human-like text, enabling seamless interactions across diverse contexts. Their responsiveness and adaptability not only foster increased user trust and a greater willingness to engage in collaborative tasks (Kirk et al., 2024), but also reduce interaction costs, thereby opening more avenues for collaboration with intelligent agents (Kim et al., 2024). Consequently, LLM agents play a critical role in sectors such as customer service, education, and healthcare, enhancing operational efficiency and enabling personalized user interactions (Kolasani, 2023; Yang et al., 2024b). However,
Seeber et al. (2020) explores AI machines functioning as teammates rather than tools. An international group of 65 collaboration scientists developed a research agenda to assess the potential risks and benefits of Machines as Teammates (MaT). This study explores the challenges associated with collaborative task completion between AI agents and humans. This highlights the research potential and necessity of investigating the dynamics of human-AI collaboration. Wiethof et al. (2021) also focus on a similar topic, and they implement a hybrid teamwork model by incorporating AI agents into the collaborative writing process, thereby addressing the research gap concerning the acceptance of AI in complementary hybrid work environments. Evaluation results indicate that the hybrid collaboration teams demonstrate potential for accepting non-human teammates. Additionally, Liao et al. (2024) conducted a randomized field experiment with 1,090 participants to examine the effect of AI agent augmentation on sales. The results showed a 5.46% increase in sales due to AI augmentation. Mediation analysis revealed that AI agents enhanced response timeliness, accuracy, and sentiment, thereby boosting sales. These findings demonstrate that AI agents improve human-AI collaboration efficiency and enhance overall organizational performance, highlighting the potential of AI agents in business success. The aforementioned studies collectively demonstrate the exceptional performance and research significance of AI agents in human-AI collaboration. Advancements in LLM technology have endowed AI agents with enhanced capabilities, enabling them to perform a variety of tasks autonomously. Consequently, this development is anticipated to foster competitive dynamics within future human-AI interaction frameworks (Kemp, 2024). Currently, research on the behavioral patterns of humans and AI agents in competitive scenarios remains insufficient.
2.2. Interaction in Competitive Environments
Competition inherently introduces greater complexity compared to cooperation, as it involves the strategic selection between cooperative and non-cooperative behaviors (Porter, 2008; Ramazi et al., 2015). While cooperative interactions are characterized by aligned interests and mutual goals, competitive settings require participants to navigate a dynamic landscape where collaboration may be pursued or withheld based on situational demands (Axelrod and Hamilton, 1981). This duality necessitates a more nuanced understanding of behavioral patterns, particularly when integrating AI agents into such environments. AI agents, traditionally perceived as supportive tools aiding users, are increasingly adopting roles that position them as autonomous external entities with their own objectives and decision-making capabilities (Bender et al., 2021). This shift from helper to competitor demands more comprehensive considerations in their interactions, as the narrative of Westworld [^2] https://westworld.fandom.com/wiki/Westworld_(TV_series) appears to possess a realistic potential for actualization.
While existing research has predominantly focused on cooperative scenarios, human–AI interactions within competitive contexts remain comparatively underexplored (Krakowski et al., 2023). Specifically, non-cooperative actions break cooperation and complicate the restoration of previous satisfaction levels. Significant effort is required to re-establish cooperation, as sustained non-cooperative behavior induces emotional fluctuations, including anger, disappointment, and distrust (Lee and Selart, 2015; Fitness, 2012). As LLM agents become increasingly integrated into society, instances of selfish and negative decision-making by LLMs are emerging (Chiang et al., 2024). However, research on human responses—emotional, behavioral, and attitudinal—to LLM agents’ non-cooperative behaviors remains insufficient.
Drawing on the behavior that emerges following non-cooperative actions in human competitive interactions, some factors have been proposed as effective measures for assessing behavioral patterns in competitive environments: benevolent actions during competition and the efforts opponents make to restore cooperation following instances of non-cooperative behavior. Initiative acts of amicability display can mitigate the adversarial nature of competition and foster more productive cooperative interactions (Fehr and Gächter, 2000b). Additionally, the propensity of competitors to forgive and attempt to re-establish cooperative relationships after experiencing non-cooperative behavior is pivotal for sustaining long-term engagement and reducing conflict (Tyler and Blader, 2003; Seppala and Cameron, 2015). These factors are instrumental in capturing the complexities of human-AI interactions in competitive settings, providing valuable insights into the underlying cooperative mechanisms. However, empirical evidence supporting these factors in the context of interactions with LLM agents remains limited. This gap underscores the need for further research to validate their effectiveness and to explore their implications for understanding how humans may modify their behavior when interacting with AI agents in competitive scenarios.
2.3. Factors in Human-AI Interaction
Extensive research has identified many factors influencing human–AI interactions. From the human perspective, these include trust (Ferrario et al., 2020), gender differences (Mou and Xu, 2017), biases towards new technology (Pratchett, 1999), and acceptance levels (Kelly et al., 2023), among others. While these factors are diverse and may lead to varied conclusions across different contexts, the gender of human participants consistently emerges as one of the most significant influences (Jang et al., 2022; Jermutus et al., 2022; Varona and Suárez, 2022). This prominence of gender is attributed not only to its direct impact on interactions with AI but also because gender often serves as a moderating variable that influences other factors, such as trust and acceptance. Consequently, gender plays a dual role in shaping the dynamics of human-AI collaboration, making it a crucial consideration for future research and the design of human-AI interaction.
@@ 2.3. Factors in Human-AI Interaction
Specially, Ahn et al. (2022) found that female participants are generally more inclined than male participants to trust and accept AI recommendations in consumer contexts. This suggests that females prefer to view AI as an assistant that helps them identify the necessary products. Additionally, Ofosu-Ampong (2023) found that gender significantly influences STEM students’ acceptance and use of AI-based tools in collaborative tasks. Female students, owing to their comparatively lower interest in technology, are more reluctant to use AI tools, limiting their opportunities to leverage AI assistance to enhance their academic performance. Renz et al. (2024) reported that females are less willing to use Strong AI than males, a disparity that is influenced by their emotional responses toward the technology. These conclusions collectively demonstrate that gender plays a significant role in human-AI interactions and collaborative processes, influencing human behavior and collaborative performance.
From the AI perspective in human-AI interaction, factors such as explainability (Arrieta et al., 2020), privacy (Curzon et al., 2021), and adaptability (Wang et al., 2023), among others, significantly influence human decision-making during these interactions. These factors can be categorized as aspects of how humans cognize the AI counterpart. This cognition is shaped by prior experiences, assumptions about the AI’s characteristics, and inherent biases, engaging in collaboration significantly impacts human attitudes and behaviors (Ishowo-Oloko et al., 2019; Burr et al., 2018). Therefore, humans’ recognition of AI’s characteristics significantly influences cooperative behavior in human-AI interactions. In other words, the transparency of AI’s characteristics plays a critical role in shaping the dynamics and outcomes of these interactions.
Liu (2021) demonstrate that machine-learning-based AI systems, characterized by lower social presence, lead to increased user uncertainty and decreased trust. However, enhancing rule transparency mitigates user uncertainty and fosters greater trust in the AI system. Notably, the mechanisms through which transparency influences trust vary across AI systems. Ishowo-Oloko et al. (2019) emphasize the trade-off between AI transparency and cooperative performance in their study. This indicates that the issue of transparency cannot be pursued in isolation; instead, specific contexts must be considered, and determining an optimal level of transparency remains an unresolved challenge. Additionally, misleading or inconsistent declarations regarding the fundamental nature of AI agents can significantly undermine user trust and negatively impact collaborative outcomes (Makovi et al., 2023). In summary, these studies collectively highlight the significant impact of AI declarations on the dynamics of human-AI collaborative interactions, shedding light on the underlying considerations regarding AI transparency, trust, and cooperative performance.
While existing research has predominantly focused on cooperative interactions between human and AI, the dynamics within competitive scenarios remain relatively underexplored. Introducing competitive settings not only complements existing studies by broadening the scope of human–AI interactions but also provides a unique framework to examine how human gender and the declarations of AI agents influence outcomes in competitive environments. Investigating these factors across cooperative and competitive contexts is essential for developing a more comprehensive understanding of human-AI interaction, offering insights with significant theoretical and practical implications.
3. Experimental Scenarios
3.1. Prisoners’ Dilemma Theory
Game theory provides a mathematical framework for analyzing multi-agent decision-making and its cooperative outcomes (Branzei et al., 2008; Rajeswaran et al., 2020). It has proven effective in computational sociology for studying and predicting human cooperative behavior (Xu et al., 2020; Jain et al., 2020). The Prisoner’s Dilemma is one of the most renowned models in game theory (Dreber et al., 2008). In this scenario, two prisoners are charged with a common offense and are interrogated separately without any communication (Crandall, 2014). Each prisoner faces a binary decision: to remain silent (cooperate) or to testify against the other (defect). This model is frequently used to analyze the dynamics of cooperation and defection, as well as the tension between pursuing self-interest and cooperating for mutual benefit (You et al., 2022; Yu, 2022).
Prior research has demonstrated that the Prisoner’s Dilemma is a highly effective model for evaluating cooperative behavior in repeated game scenarios (Bó, 2005). Psychological factors have been shown to significantly influence participants’ decision-making processes within this framework. For instance, reciprocal behaviors encourage individuals to cooperate, even without enforceable agreements, as individuals are often willing to incur personal costs to uphold fairness and trust in relationships (Fehr and Gächter, 2000a; Camerer, 2011). Emotional responses, such as empathy and anger, particularly during the reconstruction phase following the breakdown of cooperation, play a crucial role in shaping decision-making (Tversky et al., 1982). These findings underscore the effectiveness of the Prisoner’s Dilemma as a tool for exploring the psychological mechanisms underlying behaviors in competitive settings. Consequently, integrating AI agents into the Prisoner’s Dilemma framework provides a valuable opportunity to investigate human-AI interactions in competitive scenarios. This research direction remains methodologically robust and highly relevant.
3.2. Experiment Setup
Figure 1. Experimental Scenarios. (1) The overall layout of the experimental site consists of two glass rooms, where participants can face the researcher directly; (2) One of the separate rooms; (3) The selection cards for the participants and the provided record sheets.

### v2

# 111* Corresponding Author

@@ Abstract.
As large language models (LLMs) become increasingly capable of autonomous decision-making, they introduce new challenges and opportunities for human-AI cooperation in mixed-motive contexts. While prior research has primarily examined AI in assistive or cooperative roles, little is known about how humans interact with AI agents perceived as independent and strategic actors. This study investigates human cooperative attitudes and behaviors toward LLM agents by engaging 30 participants (15 males, 15 females) in repeated Prisoner’s Dilemma games with agents differing in declared identity: purported human, rule-based AI, and LLM agent. Behavioral metrics—including cooperation rate, decision latency, unsolicited cooperative acts and trust restoration tolerance—were analyzed to assess the influence of agent identity and participant gender. Results revealed significant effects of declared agent identity on most cooperation-related behaviors, along with notable gender differences in decision latency. Furthermore, qualitative responses suggest that these behavioral differences were shaped by participants’ interpretations and expectations of the agents. These findings contribute to our understanding of human adaptation in competitive cooperation with autonomous agents and underscore the importance of agent framing in shaping effective and ethical human-AI interaction.
@@ 1. Introduction
Over the past five years, the field of artificial intelligence (AI) has experienced significant growth, largely driven by the rapid advancement and deployment of large language models (LLMs) (Wu et al., 2023). At the same time, with the influx of autonomous decision-making AI agents, interactions between humans and AI agents are becoming increasingly complex. Meanwhile, existing research on human-AI interaction has primarily focused on AI as assistive tools that augment human capabilities under human supervision (Xu et al., 2023; Amershi et al., 2019; Cabrera et al., 2023). From a longer-term perspective, recent advances in LLMs have enabled AI agents to execute complex tasks autonomously—from strategic planning to conflict resolution (Luo et al., 2024). Given that more and more research underscores the effectiveness of LLMs in simulating human cognition and cooperative behavior (Ke et al., 2024; Abbasiantaeb et al., 2024), these models hold the potential to integrate into human society as quasi-autonomous partners, capable of making independent decisions during interactions with humans and thereby enhancing performance across various domains (Liu et al., 2023; Gong et al., 2025).
Despite these promising developments, human attitudes toward cooperation with LLMs remain ambiguous due to various factors. Human attitudes and behaviors toward AI agents are strongly influenced by how AI is perceived, particularly when AI transitions from a tool to an independent decision-making entity (Fahnenstich et al., 2024). This transition fundamentally changes human-AI interaction by introducing negotiation and trust-building processes traditionally reserved for human-to-human collaboration (Duan et al., 2025). Autonomous AI entities challenge norms in cooperative tasks by requiring humans to evaluate the AI’s decision-making for accuracy and alignment with human goals and values. However, when humans are faced with LLMs as autonomous partners, concerns about the “black-box” nature of these systems—that is, the opacity and difficulty in interpreting their decision-making processes (Nguyen et al., 2023; Chkirbene et al., 2024)—further exacerbate skepticism in critical decision-making scenarios.
Humans’ recognition of AI agents’ identities, along with variations in their trust attitudes towards AI, plays a crucial role in shaping cooperative decision-making. On the one hand, how AI agents present their identity—whether explicitly labeled as non-human or intentionally designed to mimic human-like traits—can significantly influence user attitudes, trust levels, and cooperative behaviors during interactions (Geiselmann et al., 2023). For example, humans tend to demonstrate greater trust and willingness to cooperate with transparent systems like rule-based AI, whereas opaque systems such as LLMs often elicit more cautious or adversarial behaviors, particularly in high-stakes collaborative scenarios (Lamparth et al., 2024). Besides, due to the perception and interaction styles, gender-specific variations in trust in AI agents have emerged as a critical factor in shaping the cooperation results in human engagement with AI agents (Jermutus et al., 2022; Varona and Suárez, 2022).
To further explore how autonomous decision-making AI agents reshape interaction dynamics, it is imperative to figure out several key factors that impact humans adapting their cooperative strategies based on the AI’s perceived agency and trust intention. Based on the research review, our study focuses on two perspectives: the declared identity of the LLM agent and the human gender. Additionally, considering the practical implications for improving cooperative interactions, we emphasize considering and deconstructing the contextual factors that affect trust dynamics, communication, and outcomes during this cooperative decision-making process. To address these issues, we employed the classic “Prisoner’s Dilemma” as an experimental framework and proposed the following research questions:
  • RQ1: How does the declared identity of the LLM agents influence human’s cooperative attitude and behavior in prisoner’s dilemmas?
  • RQ2: How does gender influence human’s cooperative attitude and behavior towards LLM agents in prisoner’s dilemmas?
  • RQ3: What other contextual factors mediate the cooperative attitude and behavior between humans and LLM agents?
This study makes four key contributions to the field of Human-AI Interaction. First, it demonstrates that the declared identity of AI agents—whether human, rule-based, or LLM—significantly affects user cooperation rates, decision latency, and trust tolerance in repeated interactions. Second, it identifies gender-related differences in decision latency, suggesting that participant gender plays a role in shaping temporal aspects of decision-making with AI partners. Third, it investigates the interaction effects between agent identity and participant gender to uncover the factors driving these behavioral disparities. Finally, it provides empirical evidence of how humans adapt their strategies in mixed-motive interactions with autonomous AI agents, advancing our understanding of interaction dynamics beyond purely cooperative paradigms.
2. Related Work
2.1. Human-AI Cooperation in the Prisoner’s Dilemma
Game theory provides a mathematical framework for analyzing decision-making scenarios where multiple agents influence outcomes, with the Prisoner’s Dilemma being a key model for studying cooperation (Branzei et al., 2008; Rajeswaran et al., 2020). Recently, game theory has been applied in computational sociology to study cooperation among agents, with a growing focus on human-AI interactions within the Prisoner’s Dilemma (Xu et al., 2020; Jain et al., 2020; Yang and Wang, 2020; Gradin et al., 2016). Oskamp (1971) and Haesevoets et al. (2018) have both highlighted that the mixed-motive scenarios presented in the prisoner’s dilemma serve as an effective measurement for examining social interaction issues within complex motivational environments.
Fontana et al. (2024) employs the Prisoner’s Dilemma to examine the cooperative tendencies and behavioral patterns of various LLMs over repeated interactions. This work introduces a systematic framework for analyzing LLM decision-making in social contexts, contributing to research on AI auditing and alignment. Ng (2023) explores human cooperative behavior with different types of partners (human, cooperative AI, non-cooperative AI) through the framework of the Prisoner’s Dilemma. This work highlights the value of the Prisoner’s Dilemma in studying human-AI collaboration dynamics.
In the study of human behavior within the Prisoner’s Dilemma, numerous variables have been proposed as indicators to assess interaction, collectively offering valuable insights into how individuals navigate strategic dilemmas. These indicators, ranging from cooperation rates to reaction times, reveal the interplay between rational deliberation and social strategies in shaping human decision-making. For example, Wang et al. (2020) identified decision latency in iterative Prisoner’s Dilemma games as a key metric of deliberative processes, reflecting the extent to which participants carefully consider their opponent’s strategy, which ultimately influences the overall rate of cooperation. Additionally, Wang et al. (2020) introduced unconditional cooperation—defined as unsolicited cooperative acts—as another behavioral measure. They found that individuals are inclined to reciprocate proactive cooperation and often make attempts to establish mutual cooperation before unilateral defection occurs, highlighting the importance of proactive strategies in fostering cooperation. Furthermore, forgiveness strategies offer an effective means of restoring cooperative relationships. Specifically, Martinez-Vaquero et al. (2015) demonstrated that forgiving an opponent’s betrayal enables participants to achieve higher payoffs over the course of long-term interactions. Moreover, the effort expended by opponents to seek forgiveness after a betrayal serves as a valuable metric for assessing behavioral decision-making within the context of the Prisoner’s Dilemma. Together, these studies illustrate that decision-making in the Prisoner’s Dilemma is not only influenced by rational calculation but also shaped by emotional and social dynamics inherent in human interactions.
As AI agents increasingly function as independent decision-makers, their behavior closely aligns with the strategic dynamics of the Prisoner’s Dilemma. This makes the model particularly suitable for exploring cooperation, trust, and forgiveness in human-AI relationships.
2.2. Effects of AI Identity on Human Cooperation Dynamics
2.2.1. AI Identity Transparency Enhances Human Trust and Cooperation
@@
As AI technology proliferates across industries, human-AI interactions are becoming increasingly complex with the influx of autonomous decision-making agents. Large Language Models (LLMs), representing a significant advancement in AI, enable LLM-driven agents to comprehend and generate human-like text, facilitating seamless interactions across diverse contexts (Chang et al., 2024). Characterized by rapid response times and high adaptability, these models foster increased user trust and willingness to engage in collaborative tasks (Kirk et al., 2024). By reducing interaction costs, they thereby expand avenues for human collaboration with intelligent agents (Kim et al., 2024; Liao et al., 2024).
While LLM-driven agents demonstrate remarkable capabilities in human-like communication, a critical aspect of their deployment is how their identity as AI agents is disclosed or represented to users. The way AI agents present their identity—whether explicitly labeled as non-human or intentionally designed to mimic human-like traits—can significantly influence user attitudes, trust levels, and cooperative behaviors during interactions (Geiselmann et al., 2023). Previous studies suggest that when AI agents transparently disclose their identity, users may exhibit higher levels of trust and willingness to collaborate, as transparency fosters confidence in the agent’s intentions (Saffarizadeh et al., 2024). Additionally, Von Eschenbach (2021) proved that transparency is a necessary condition for trust and eventually for judging AI to be trustworthy by philosophical analysis of trust. Conversely, when AI agents without disclosing their non-human identity, users may initially demonstrate high engagement (Chen et al., 2025). However, such interactions risk undermining trust if users later perceive the agent’s behavior as deceptive or misleading, highlighting the potential ethical and practical challenges in designing AI systems with varying levels of identity transparency (Li et al., 2024).
Despite these insights, current research primarily focuses on AI agents in general, without systematically addressing how different types of AI agents—such as purported real human, rule-based algorithms, and LLM-driven agents—interact with users under varying identity disclosure conditions. This lack of differentiation overlooks the possibility that user perceptions, attitudes, and trust dynamics may vary significantly across these distinct types of AI agents.
2.2.2. Human Cooperative Attitude differs between rule-based systems and LLMs
Existing research highlights that human attitude and cooperation vary significantly depending on AI characteristics such as interpretability and transparency (Wang et al., 2023). Humans may demonstrate greater trust and willingness to cooperate with transparent systems like rule-based AI. In contrast, opaque systems such as LLMs often elicit more cautious or adversarial behaviors, especially in high-stakes collaborative scenarios (Lamparth et al., 2024). For instance, rule-based systems are often preferred for their predictability and interpretability, fostering trust in contexts requiring accountability (Rane et al., 2023). Conversely, LLMs, despite their human-like communication abilities, are perceived as “black-box” systems due to their lack of transparency (Chkirbene et al., 2024), which can lead to skepticism in critical decision-making scenarios.
These differences in system design influence trust and cooperative behavior: rule-based systems can facilitate stable and predictable collaboration, whereas LLMs can foster emotional connection while introducing uncertainty (Lu et al., 2024). For instance, Panagoulias et al. (2023) emphasizes the lack of trust in LLMs in complex and high-stakes decision-making scenarios, such as in healthcare. He suggests that combining rule-based systems with LLMs can enhance their credibility. This also suggests that while LLMs excel in generalization, they may disrupt cooperation in applications requiring high transparency or accountability. These differences influence user attitudes and cooperation behavior across different contexts, reflecting the importance of AI type in shaping human behavior.
In summary, existing research primarily examines the applicability of different AI types across various contexts, often focusing on trust and usability issues while treating AI as tools or passive assistants. This perspective fails to address how AI agents that make autonomous decisions shape cooperation dynamics. Such agents introduce nuanced patterns that require humans to adapt their cooperative strategies based on the AI’s perceived agency, which are critical for optimizing human-AI collaboration, as they directly impact decision outcomes and negotiation processes.
2.3. Human Gender Matters in Cooperation Attitudes with AI Agents
While existing research has established the effects of AI identity representation and system types on human cooperation(Ferrario et al., 2020; Pratchett, 1999; Kelly et al., 2023; Xu et al., 2025), individual differences like gender remain underexplored in shaping these dynamics. Especially, gender-specific variations in trust, perception, and interaction styles have emerged as critical factors in shaping how humans engage with AI agents, highlighting the need for a deeper understanding of these dynamics (Jermutus et al., 2022; Varona and Suárez, 2022).
@@ 2.3. Human Gender Matters in Cooperation Attitudes with AI Agents
Ahn et al. (2022) found that female participants are generally more inclined than male participants to trust and accept AI recommendations in consumer contexts. This suggests that females prefer to view AI as an assistant that helps them identify the necessary products. At the same time, Ofosu-Ampong (2023) found that gender significantly influences STEM students’ acceptance and use of AI-based tools in collaborative tasks. Due to their comparatively lower interest in technology, female students are more reluctant to use AI tools, limiting their opportunities to leverage AI assistance to enhance their academic performance. Renz et al. (2024) reported that females are less willing to use Strong AI than males, a disparity that is influenced by their emotional responses toward the technology.
These conclusions collectively demonstrate that which gender the participant belongs to may play a significant role in human-AI interactions and collaborative processes because of the emotional responses and influencing human behavior and collaborative performance.
In summary, prior studies highlight the significance of gender in shaping attitudes toward and acceptance of AI systems, which concentrated on individual preferences or decision-making contexts, leaving the dynamic nature of human-AI collaboration largely unexamined.
3. Method
To address the research questions outlined above, we experimented to compare the cooperation attitude and behavior of humans towards LLM agents with different declared identities. Based on the specific LLM agents’ declared identities and related settings, we conducted the formal experiment enrolling 30 participants after a pilot study for exploration. Within the framework of our experimental design, two independent variables were carefully defined: the declared identity of the AI agent, which included purported human, purported rule-based AI agents, and LLM-based AI agents, and the gender of the participants, categorized as male or female.
3.1. Pre-experimental Setting
3.1.1. Declared Identity Setting